A chain hangs over a nail with 2.0 m on one side and 6.0 m on the other side. If the

force of friction is equal to the weight of 1.0 m of the chain, calculate the time required

for the chain to slide off the nail.
1
Expert's answer
2021-10-31T18:10:56-0400

"\lambda=m/L"


Equation of motion


"\lambda y(t)g-\lambda(L-y(t))g-\lambda g=L\lambda\frac{d^2y(t)}{dt^2}\to"


"y(t)=\frac{L+1}{2}+Ae^{\sqrt{2g/L}\cdot t}"


"L=8\ m"

"y(0)=6\ m" . So, we have "A=1.5"


"y(t)=\frac{L+1}{2}+1.5e^{\sqrt{2g/L}\cdot t}=4.5+1.5e^{1.581t}"


For "y(t)=8\ m\to t\approx0.54\ (s)" . Answer